A business retains 79% of its customers from one year to the next. If the business uses a 11% discount rate, calculate the margin multiple. Rounding: 2 decimal places.

Show work
Answer #1

Margin multiple = r/(1+i-r)

r = retention rate

i = discount rate

Margin multiple = 0.79 /(1+0.11-0.79) = 2.47